I HAVE THE PROBLEM TO GRAPH Y-1=-(x-3)
I don't understand how to use the distribtive property with just an - and no number

Answer by fcabanski(1391)

- is the same as -1.


y-1 =-(x-3) = y-1 =-x + 3


y = -x + 4